Sheffield Corp. reported the following year-end information: Beginning work in process $56000 inventory Beginning raw materiaCrane Company reported the following year-end information: Beginning work in process $32000 inventory Beginning raw materialsUsing the following information, compute the direct materials used. Raw materials inventory, January 1 $10000 Raw materials iAssuming that the direct materials used are $1900000, compute the total manufacturing costs using the following information.
Answer #1

Part 1)

The correct answer is $ 1408000

Calculation

Cost of goods manufactured

= opening wip + direct material + direct labour + manufacturing overheads- ending wip

= 56000 + (33000+850000-17000)+ 440000 + 100000 - 54000

=$ 1408000

Part 2)

The correct answer is $ 924000

Formula

Total cost of work in process = beginning wip + total manufacturing costs

= 32000 + ( 14000+ 560000 - 12000) + 210000 + 120000

= $ 924000

Part 3)

The correct answer is $ 1580000

Direct material used = opening raw material inventory + purchase - ending inventory

= 10000 + 1600000 - 30000

= $ 1580000

Part 4)

The correct answer is $ 3260000

Total manufacturing cost

= direct material used + direct labour +factory utilities + indirect labour ++ factory depreciation

= 1900000+ 760000 + 150000+ 50000+400000

= $ 3260000